OpenStudy (anonymous):

If you are riding on a Ferris wheel that is 60 feet tall and it takes you 5 minutes to make one revolution, what is your linear velocity?

OpenStudy (anonymous):

v = x / t. The x variable is distance traveled, and t is the time it took to travel the distance x. The v variable is the linear velocity. V= 60/5 = 12